OpenStudy (anonymous):

GCF of -10c^2d and 15cd^2

OpenStudy (anonymous):

5cd

OpenStudy (kc_kennylau):

1. let's break -10c^2d into its components 2. let's break 15cd^2 into its components

OpenStudy (kc_kennylau):

@sourwing please don't just give out the answer

OpenStudy (anonymous):

I suck at explaining simple algebra lol

OpenStudy (anonymous):

well that's what i thought it was but you dont add the like ^2 so it's like writing it out like 10 is 5*2 and 15 is 5*3 and so on right?

OpenStudy (kc_kennylau):

yep, and c^2 is c*c

OpenStudy (jdoe0001):

\(\large \begin{array}{ccccll} &-10c^2d &\quad and \quad &15cd^2\\ &\uparrow&&\uparrow\\ &-2\cdot \color{blue}{5}\cdot \color{blue}{c}\cdot c\cdot \color{blue}{d}&&3\cdot \color{blue}{5}\cdot \color{blue}{c}\cdot \color{blue}{d}\cdot d \end{array}\) what do you think?

OpenStudy (anonymous):

woo!

OpenStudy (kc_kennylau):

@jdoe0001 wow \(\LaTeX\) pro :D

OpenStudy (jdoe0001):

hehe, cheating using a latex editor =)